# Med-2D SegNet: A Light Weight Deep Neural Network for Medical 2D Image Segmentation

### v1

@@ 1 Introduction
Advancements in automated segmentation have transformed tasks such as tumor detection in MRI [22] and organ segmentation in CT scans [31], demonstrating how machine learning models now tackle increasingly complex challenges with remarkable accuracy [40, 30, 33, 35, 39]. Traditional CNN-based architectures like U-Net [41] and DeepLab [13] are well-regarded for their ability to extract detailed features with a reasonable computational cost, achieving widespread adoption in medical image analysis. However, the emergence of Transformer-based models has marked a paradigm shift, particularly with Vision Transformers (ViT) [44], which excel at segmenting high-resolution images. Despite their improved accuracy, Transformers demand substantial computational resources due to their higher parameter counts, making their deployment in real-world medical imaging systems challenging. This juxtaposition between CNNs and Transformers underscores a pivotal question in segmentation research: how to leverage advanced architectures while addressing their scalability and resource-intensiveness for practical applications. Segmentation models for multi-domain tasks often struggle with a trade-off between accuracy and computational cost. Models with large parameter counts excel in handling diverse applications but incur high resource requirements [37], whereas smaller models are frequently limited to single-domain tasks [18]. To address this limitation, we proposed a lightweight segmentation model that bridges the gap between parameter efficiency and multi-domain accuracy. The proposed architecture reduces the number of parameters and weights while maintaining competitive performance across domains. Moreover, the model’s ability to adapt to varying batch sizes challenges the common notion that segmentation models are inherently batch size dependent. This adaptability not only enhances its robustness but also demonstrates its suitability for medical applications, where precision and computational efficiency are critical. By offering rapid inference without sacrificing accuracy, the proposed model balances cutting-edge performance with practical utility.
@@ 2 Related Work
Recent advancements in medical image segmentation have aimed to address the limitations of task-specific models, often designed for particular imaging modalities such as computed tomography (CT) or magnetic resonance imaging (MRI), and specific anatomical regions, including organs or tumors [12, 20, 25]. The Hermes model represents a significant innovation, providing a universal framework capable of tackling diverse segmentation tasks across various body parts and imaging techniques. Inspired by the training methodologies of radiology residents, Hermes employs context-prior learning to integrate task-specific and scan-specific information, enhancing its understanding of interconnections between tasks. Its exceptional performance across eleven datasets, including CT, positron emission tomography (PET), and MRI, underscores its dynamism and proficiency in managing novel challenges [32, 19].
In parallel, the Segment Anything Model (SAM) introduced a groundbreaking approach by enabling interactive capabilities through user prompts, facilitating the segmentation of any object, including those in medical images[28]. SAM’s impressive performance across various benchmarks highlights its potential for robust segmentation in clinical settings, positioning it as a valuable complement to models like Hermes [19].
Additionally, M2SNET proposes an innovative framework to enhance segmentation accuracy by addressing limitations in existing U-shaped architectures, such as U-Net [49, 41]. Traditional methods rely on operations like element-wise addition or concatenation, which can introduce redundancy. In contrast, M2SNET employs a subtraction unit (SU) to extract unique information from adjacent feature levels, improving localization accuracy [49]. It also incorporates a multi-scale approach and a LossNet module to refine the segmentation process and enhance convergence during training.
@@ 3.1 Med Block
The Med Block utilizes a combination of feature map expansion and convolution to effectively capture multi-scale features. Initially, the input feature map $\alpha$ is expanded by a factor of 6, enhancing the model’s capacity to process a more diverse set of features. This expansion is expressed as $\tilde{\alpha}(i,j,\gamma)=6\cdot\alpha(i,j,\gamma)$ where $(i,j)$ are the spatial dimensions of the feature map, and $\gamma$ denotes the output feature maps. The expanded feature map $\tilde{\alpha}$ is then passed through a $1\times 1$ convolution operation, denoted by $\Psi(1\times 1,\tilde{\alpha},\gamma)$, to produce the refined feature map $\Omega_{ex}(i,j,\gamma)=\Psi(1\times 1,6\cdot\alpha(i,j,\gamma))$The entire operation of the layer, combining both expansion and convolution, can be represented as:
@@
$=\sum_{\alpha=1}^{C}\Psi(1\times 1,6\cdot\alpha(i,j,\gamma))\quad\text{for }\gamma\in\{1,2,\ldots,6C\}$
@@
This equation describes a $7\times 7$ convolution operation, where $p-4$ and $q-4$ center the filter over the pixel $(i,j)$, ensuring that the filter symmetrically captures neighboring pixels within the range from $(i-3,j-3)$ to $(i+3,j+3)$. The element-wise multiplication $\Phi(i+p-4,j+q-4,c)\cdot\Psi_{c}(p,q)$ allows the model to learn localized spatial features, effectively capturing patterns and spatial dependencies.
@@
for $\text{for }\gamma\in\left\{1,2,\ldots,\frac{F_{1}}{24}\right\}.$
@@ 4 Parameter Complexity in the Med Block Architecture
for $\text{for }\gamma\in\left\{1,2,\ldots,\frac{F_{1}}{24}\right\}.$
@@ 5.1 Dataset and Preprocessing
To thoroughly assess the effectiveness and resilience of our model, we conducted an extensive series of experiments across 20 publicly available datasets which are represented in Table 1, carefully selected for their diversity in modality, anatomical regions, and clinical targets. Notably, some datasets include 3D MRI images, which we converted into 2D slices through preprocessing steps to align with our approach. To enhance the generalization ability of the model, we applied a wide range of data augmentation techniques [10] and fine-tuned hyperparameters. Each dataset was divided into 80% for training, 10% for validation, and 10% for testing. In certain cases, to ensure consistency and fairness in comparison, we adhered to the same dataset split methods as prior works; for example, DUCK-Net [16] for polyp segmentation and U-Mamba [33] for EndoVis17. This careful design ensures that our evaluation reflects the true resilience of our model across a wide spectrum of biomedical imaging challenges. More comparisons can be found in the supplementary material Experimental Setup section.
@@ 6.1 Experimental Results
Our proposed method exhibits remarkable consistency and adaptability across prominent polyp segmentation benchmarks, achieving competitive performance that underscores its robustness in diverse clinical scenarios. As displayed in Table 1, the model attained outstanding Dice scores 0.9578 on Kvasir-SEG [26], 0.9494 on Kvasir-Capsule [43], 0.9605 on CVC-ClinicDB [5], 0.9478 on CVC-ColonDB [46], and 0.9538 on ETIS-LARIBPOLYPDB [6]. These consistently high scores reveal the model’s ability to handle the nuanced variations inherent to each dataset.
@@
Polyp Kvasir-SEG [26] 256 x 256 0.9578
Kvasir-Capsule [43] 256 x 256 0.9494
CVC-ClinicDB [5] 256 x 256 0.9605
CVC-ColonDB [46] 256 x 256 0.9478
ETIS-LARIBPOLYPDB [6] 256 x 256 0.9538
Dermoscopy ISIC 2018 [45] 256 x 256 0.9128
@@
Fundus DRIVE [3] 512 x 512 0.8182
CHASE-DB1 [17] 512 x 512 0.7767
Retinography RIM One [4] 128 x 128 0.8834
Ophthalmology Ravir [23] 512 x 512 0.7705
X-Rays Chest X-Ray (Lung) [11] 256 x 256 0.9784
Ultrasound Breast Ultrasound Images [1] 256 x 256 0.7997
Histology CoNIC [21] 256 x 256 0.8356
GLaS (Test A) [42] 256 x 256 0.9125
GLaS (Test B) [42] 256 x 256 0.9056
HuBMAP (Kidney) [24] 256 x 256 0.8338
MRI Brain Tumor [9] 256 x 256 0.9144
CT LiTs [8] 256 x 256 0.9202
Endoscopy Endovis 2017 (Binary) [2] 256 x 256 0.9307
Endovis 2017 (Multiclass) [2] 256 x 256 0.7422
Hematology Cell Nuclic Segmentation (Blood) [15] 512 x 512 0.9267
@@
PVT-EMCAD-B2 [40] 26.76 0.9275
DUCK-Net (34 filters) [16] 155.4 0.9502
RAPUNet [30] 43.85 0.9390
Kvasir-SEG [26] Ours 2.07 0.9578
PVT-EMCAD-B2 [40] 26.76 0.9521
DUCK-Net (34 filters) [16] 155.4 0.9478
RAPUNet [30] 43.85 0.9610
CVC-ClinicDB [5] Ours 2.07 0.9605
PVT-EMCAD-B2 [40] 26.76 0.9231
DUCK-Net (17 filters) [16] 38.92 0.9353
RAPUNet [30] 43.85 0.9526
CVC-ColonDB [46] Ours 2.07 0.9478
PVT-EMCAD-B2 [40] 26.76 0.9229
DUCK-Net (34 filters) [16] 155.4 0.9354
RAPUNet [30] 43.85 0.9572
@@
PVT-GCASCADE [39] 3.32 0.9151
TransFuse [48] 26.3 0.9010
UCTransNet [47] 65.6 0.9050
ISIC 2018 [14] Ours 2.07 0.9128
@@
DRIVE [3] Ours 2.07 0.8182
nnU-Net [25] 33 0.6264
nnFormer [50] 60 0.6135
U-Mamba_Bot [35] 63 0.6540
Swin-UMamba [33] 28 0.6783
Endovis [2] Ours 2.07 0.7422
@@ 7 Exploring Zero-shot Learning
The results, presented in Table 3, illustrate the performance of our zero-shot learning model compared to the established ResUNet++ framework, both with and without Test-Time Augmentation (TTA) [27], using the DSC as a key performance metric. Notably, our model achieved a DSC score of 0.8319 on the CVC-ClinicDB dataset when trained on Kvasir-Seg, significantly surpassing ResUNet++ (0.6468), ResUNet++ with TTA (0.6737) and RAPUNet (0.8214). Additionally, our model demonstrated strong zero-shot segmentation capabilities across other datasets, attaining DSC scores of 0.9397 on the PH2 dataset (trained on ISIC 2018), 0.7737 on ISIC 2018 (trained on PH2), and 0.7386 on CHASE-DB1 (trained on DRIVE).

### v2

@@ 1 Introduction
Advancements in automated segmentation have transformed tasks such as tumor detection in MRI [gtifa2024multimodal] and organ segmentation in CT scans [li2023eres], demonstrating how machine learning models now tackle increasingly complex challenges with remarkable accuracy [rahman2024emcad, lee2024metaformer, liu2024swin, ma2024u, rahman2024g]. Traditional CNN-based architectures like U-Net [ronneberger2015u] and DeepLab [chen2017deeplab] are well-regarded for their ability to extract detailed features with a reasonable computational cost, achieving widespread adoption in medical image analysis. However, the emergence of Transformer-based models has marked a paradigm shift, particularly with Vision Transformers (ViT) [thisanke2023semantic], which excel at segmenting high-resolution images. Despite their improved accuracy, Transformers demand substantial computational resources due to their higher parameter counts, making their deployment in real-world medical imaging systems challenging. This juxtaposition between CNNs and Transformers underscores a pivotal question in segmentation research: how to leverage advanced architectures while addressing their scalability and resource-intensiveness for practical applications. Segmentation models for multi-domain tasks often struggle with a trade-off between accuracy and computational cost. Models with large parameter counts excel in handling diverse applications but incur high resource requirements [nguyen2021multi], whereas smaller models are frequently limited to single-domain tasks [galdran2022state]. To address this limitation, we proposed a lightweight segmentation model that bridges the gap between parameter efficiency and multi-domain accuracy. The proposed architecture reduces the number of parameters and weights while maintaining competitive performance across domains. Moreover, the model’s ability to adapt to varying batch sizes challenges the common notion that segmentation models are inherently batch size dependent. This adaptability not only enhances its robustness but also demonstrates its suitability for medical applications, where precision and computational efficiency are critical. By offering rapid inference without sacrificing accuracy, the proposed model balances cutting-edge performance with practical utility.
@@ 2 Related Work
Recent advancements in medical image segmentation have primarily focused on overcoming the limitations inherent to modality-specific and organ-specific models. Traditional methods such as U-Net have provided foundational architectures for segmentation tasks but face constraints like redundancy in feature integration and limited adaptability to diverse medical imaging modalities [ronneberger2015u, isensee2021nnu]. Addressing these limitations, ConvNeXt [liu2022convnet] introduced an inverse bottleneck block, achieving significant improvements in vision tasks by emphasizing efficient feature extraction and computational efficiency. MedNeXt [roy2023mednext] adapted this concept effectively to 3D medical imaging, demonstrating superior performance in volumetric segmentation tasks by capturing rich spatial context.
Further advancements in 2D medical image segmentation have seen the development of architectures such as TransFuse [zhang2021transfuse] and Swin-Unet [cao2023swin]. TransFuse merges CNN and Transformer modules to leverage both local spatial details and global context, resulting in enhanced segmentation accuracy. Similarly, Swin-Unet employs a Swin Transformer backbone to efficiently capture hierarchical representations, significantly improving model generalization across various medical image segmentation tasks.
Additionally, recent architectures like M2SNET [zhao2023m] and M2U-Net [laibacher2019m2u] propose innovative approaches addressing specific shortcomings of traditional U-shaped frameworks. M2SNET introduces a subtraction unit (SU) designed to extract unique information between adjacent feature levels, effectively reducing redundancy and enhancing localization accuracy. It further integrates multi-scale processing and LossNet modules to refine segmentation quality and convergence during training. M2U-Net, on the other hand, adopts multi-level features aggregation to efficiently integrate contextual information, improving segmentation accuracy across different medical imaging modalities. These advancements highlight ongoing efforts toward developing versatile, accurate, and computationally efficient segmentation frameworks capable of effectively handling diverse medical imaging datasets and tasks.
@@ 3.1 Med Block
The Med Block enhances multi-scale feature extraction by expanding the feature map dimensionality, not merely scaling its values. Specifically, given an input feature map $\alpha\in\mathbb{R}^{H\times W\times C}$, a $1\times 1$ convolution is applied to project it into a higher-dimensional space, resulting in an expanded feature map $\tilde{\alpha}\in\mathbb{R}^{H\times W\times 6C}$. This expansion enriches the representational capacity of the network by increasing the number of channels sixfold, thereby enabling the model to capture more diverse and complex features without altering the spatial resolution. The entire operation of the layer, combining both expansion and convolution, can be represented as:
@@
$\sum_{\alpha=1}^{C}\Psi(1\times 1,6\cdot\alpha(i,j,\gamma))\quad\text{for }\gamma\in\{1,2,\ldots,6C\}$
@@
This equation describes a $7\times 7$ convolution operation, where p4p-4 and q4q-4 center the filter over the pixel $(i,j)$, ensuring that the filter symmetrically captures neighboring pixels within the range from (i3,j3)(i-3,j-3) to $(i+3,j+3)$. The element-wise multiplication $\Phi(i+p-4,j+q-4,c)\cdot\Psi_{c}(p,q)$ allows the model to learn localized spatial features, effectively capturing patterns and spatial dependencies.
@@
$\text{for }\gamma\in\left\{1,2,\ldots,\frac{F_{1}}{24}\right\}.$
@@ 4 Parameter Complexity in the Med Block Architecture
$\text{for }\gamma\in\left\{1,2,\ldots,\frac{F_{1}}{24}\right\}.$
@@ 5.1 Dataset and Preprocessing
To thoroughly assess the effectiveness and resilience of our model, we conducted an extensive series of experiments across 20 publicly available datasets which are represented in Table 1, carefully selected for their diversity in modality, anatomical regions, and clinical targets. Notably, some datasets include 3D MRI images, which we converted into 2D slices through preprocessing steps to align with our approach. To enhance the generalization ability of the model, we applied a wide range of data augmentation techniques [buslaev2020albumentations] and fine-tuned hyperparameters. Each dataset was divided into 80% for training, 10% for validation, and 10% for testing. In certain cases, to ensure consistency and fairness in comparison, we adhered to the same dataset split methods as prior works; for example, DUCK-Net [dumitru2023using] for polyp segmentation and U-Mamba [liu2024swin] for EndoVis17. This careful design ensures that our evaluation reflects the true resilience of our model across a wide spectrum of biomedical imaging challenges. More comparisons can be found in the supplementary material Experimental Setup section.
@@ 6.1 Experimental Results
Our proposed method exhibits remarkable consistency and adaptability across prominent polyp segmentation benchmarks, achieving competitive performance that underscores its robustness in diverse clinical scenarios. As displayed in Table 1, the model attained outstanding Dice scores 0.9578 on Kvasir-SEG [jha2020kvasir], 0.9494 on Kvasir-Capsule [smedsrud2021kvasir], 0.9605 on CVC-ClinicDB [bernal2015wm], 0.9478 on CVC-ColonDB [vazquez2017benchmark], and 0.9538 on ETIS-LARIBPOLYPDB [bernal2017comparative]. These consistently high scores reveal the model’s ability to handle the nuanced variations inherent to each dataset.
@@
Polyp Kvasir-SEG [jha2020kvasir] 256 x 256 0.9578
Kvasir-Capsule [smedsrud2021kvasir] 256 x 256 0.9494
CVC-ClinicDB [bernal2015wm] 256 x 256 0.9605
CVC-ColonDB [vazquez2017benchmark] 256 x 256 0.9478
ETIS-LARIBPOLYPDB [bernal2017comparative] 256 x 256 0.9538
Dermoscopy ISIC 2018 [tschandl2018ham10000] 256 x 256 0.9128
@@
Fundus DRIVE [asad2014ant] 512 x 512 0.8182
CHASE-DB1 [fraz2012ensemble] 512 x 512 0.7767
Retinography RIM One [batista2020rim] 128 x 128 0.8834
Ophthalmology Ravir [hatamizadeh2022ravir] 512 x 512 0.7705
X-Rays Chest X-Ray (Lung) [candemir2013lung] 256 x 256 0.9784
Ultrasound Breast Ultrasound Images [ALDHABYANI2020104863] 256 x 256 0.7997
Histology CoNIC [graham2024conic] 256 x 256 0.8356
GLaS (Test A) [sirinukunwattana2017gland] 256 x 256 0.9125
GLaS (Test B) [sirinukunwattana2017gland] 256 x 256 0.9056
HuBMAP (Kidney) [hubmap-hacking-the-human-vasculature] 256 x 256 0.8338
MRI Brain Tumor [Buda_2019] 256 x 256 0.9144
CT LiTs [bilic2023liver] 256 x 256 0.9202
Endoscopy Endovis 2017 (Binary) [allan20192017] 256 x 256 0.9307
Endovis 2017 (Multiclass) [allan20192017] 256 x 256 0.7422
Hematology Cell Nuclic Segmentation (Blood) [DEPTO2021101653] 512 x 512 0.9267
@@
PVT-EMCAD-B2 [rahman2024emcad] 26.76 0.9275
DUCK-Net (34 filters) [dumitru2023using] 155.4 0.9502
RAPUNet [lee2024metaformer] 43.85 0.9390
Kvasir-SEG [jha2020kvasir] Ours 2.07 0.9578
PVT-EMCAD-B2 [rahman2024emcad] 26.76 0.9521
DUCK-Net (34 filters) [dumitru2023using] 155.4 0.9478
RAPUNet [lee2024metaformer] 43.85 0.9610
CVC-ClinicDB [bernal2015wm] Ours 2.07 0.9605
PVT-EMCAD-B2 [rahman2024emcad] 26.76 0.9231
DUCK-Net (17 filters) [dumitru2023using] 38.92 0.9353
RAPUNet [lee2024metaformer] 43.85 0.9526
CVC-ColonDB [vazquez2017benchmark] Ours 2.07 0.9478
PVT-EMCAD-B2 [rahman2024emcad] 26.76 0.9229
DUCK-Net (34 filters) [dumitru2023using] 155.4 0.9354
RAPUNet [lee2024metaformer] 43.85 0.9572
@@
PVT-GCASCADE [rahman2024g] 3.32 0.9151
TransFuse [zhang2021transfuse] 26.3 0.9010
UCTransNet [UCTransNet] 65.6 0.9050
ISIC 2018 [codella2019skin] Ours 2.07 0.9128
@@
DRIVE [asad2014ant] Ours 2.07 0.8182
nnU-Net [isensee2021nnu] 33 0.6264
nnFormer [zhou2023nnformer] 60 0.6135
U-Mamba_Bot [ma2024u] 63 0.6540
Swin-UMamba [liu2024swin] 28 0.6783
Endovis [allan20192017] Ours 2.07 0.7422
@@ 7 Exploring Zero-shot Learning
The results, presented in Table 3, illustrate the performance of our zero-shot learning model compared to the established ResUNet++ framework, both with and without Test-Time Augmentation (TTA) [jha2021comprehensive], using the DSC as a key performance metric. Notably, our model achieved a DSC score of 0.8319 on the CVC-ClinicDB dataset when trained on Kvasir-Seg, significantly surpassing ResUNet++ (0.6468), ResUNet++ with TTA (0.6737) and RAPUNet (0.8214). Additionally, our model demonstrated strong zero-shot segmentation capabilities across other datasets, attaining DSC scores of 0.9397 on the PH2 dataset (trained on ISIC 2018), 0.7737 on ISIC 2018 (trained on PH2), and 0.7386 on CHASE-DB1 (trained on DRIVE).